# CaKE: Circuit-aware Editing Enables Generalizable Knowledge Learners

### v1

@@ Abstract
Knowledge Editing (KE) enables the modification of outdated or incorrect information in large language models (LLMs). While existing KE methods can update isolated facts, they struggle to generalize these updates to multi-hop reasoning tasks that depend on the modified knowledge. Through an analysis of reasoning circuits—the neural pathways LLMs use for knowledge-based inference, we observe that current layer-localized KE approaches, such as MEMIT and WISE, which edit only single or a few model layers, struggle to effectively incorporate updated information into these reasoning pathways. To address this limitation, we propose CaKE (Circuit-aware Knowledge Editing), a novel method that enables more effective integration of updated knowledge in LLMs. CaKE leverages strategically curated data, guided by our circuits-based analysis, that enforces the model to utilize the modified knowledge, stimulating the model to develop appropriate reasoning circuits for newly integrated knowledge. Experimental results show that CaKE enables more accurate and consistent use of updated knowledge across related reasoning tasks, leading to an average of 20% improvement in multi-hop reasoning accuracy on MQuAKE dataset compared to existing KE methods. We release the code and data in https://github.com/zjunlp/CaKE.
@@
Yunzhi ${}^{\text{\Scorpio\Pisces}}$thanks: Work done during Yunzhi’s visit to UCLA., Jizhan ${}^{\text{\Scorpio}}$, Jia-Chen ${}^{\text{\Pisces}}$, Ningyu ${}^{\text{\Scorpio}}$22footnotemark: 2, Shumin ${}^{\text{\Cancer}}$, Huajun ${}^{\text{\Scorpio}}$thanks: Corresponding Authors, Nanyun ${}^{\text{\Pisces}}$22footnotemark: 2, ${}^{\text{\Scorpio}}$ Zhejiang University ${}^{\text{\Cancer}}$ National University of Singapore ${}^{\text{\Pisces}}$ University of California, Los Angeles {yyztodd,fangjizhan,zhangningyu,huajunsir}@zju.edu.cn gujc@ucla.edu,violetpeng@cs.ucla.edu,shumin@nus.edu.sg
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in diverse tasks Yang et al. (2024a); Dubey et al. (2024); OpenAI (2024); Guo et al. (2025), achieving performance that rivals or even exceeds human experts. However, their practical deployment faces some critical limitations: parametric knowledge remains static after pretraining, making it challenging to keep up with evolving real-world information; their propensity for hallucinations also undermines reliability. Knowledge editing (KE) has emerged as a promising solution to update the knowledge in models precisely Mitchell et al. (2021); Wang et al. (2024c); Jiang et al. (2025). Although existing KE methods achieve good results on simple factual updates Yao et al. (2023); Zhang et al. (2024b), they often exhibit fundamental limitations: edits propagate inconsistently through related knowledge structures and downstream reasoning tasks Cohen et al. (2024); Qin et al. (2024); Yao et al. (2023); excessive focus on surface-level pattern matching Hoelscher-Obermaier et al. (2023), and locality issues for other unrelated knowledge and general ability Gu et al. (2024); Gupta et al. (2024).
@@
Our work specifically addresses the poor performance of edited models in downstream reasoning tasks that involve the updated knowledge Zhong et al. (2023); Zhang et al. (2024d). Consider a representative case in Figure 1 : after editing ‘Eddie Mathews, citizenship, United States → United Kingdom’, models correctly answer direct queries but fail multi-hop reasoning like ‘The capital of the country that Eddie Mathews was a citizen of is?’ (still outputting ‘Washington D.C.’). Critically, this is not merely an editing artifact: vanilla LLMs often correctly answer single-hop questions while failing their multi-hop counterparts Yang et al. (2024c); Biran et al. (2024), suggesting deeper architectural limitations in knowledge utilization.
@@
We trace these limitations to a misalignment between KE strategies and the inherent reasoning architectures of LLMs. To investigate this disconnect, we examine how LLMs leverage knowledge in downstream reasoning tasks. Recent analysis suggests that knowledge is not merely statically stored but dynamically activated through specialized circuits Yao et al. (2024); Biran et al. (2024); Yu et al. (2025). However, these analyses overlook the phenomenon of LLM failures in reasoning circuits and fail to explore the underlying causes. Our investigation (§2) delves deeper into reasoning circuits, analyzing their structure and identifying the reasons behind failures in multi-hop reasoning cases. Specifically, the multi-hop reasoning emerges from coordinated computing circuits: early layers handle the first hop, extracting the bridge entity at the end-token of the first hop. This bridge entity, along with second-hop relation information, is then routed to the last token position in the middle layers. Subsequently, later layers utilize this information at the last token position to complete the reasoning process (Figure 2 (a)). We then analyze the entity and relation information at the last token position in failed multi-hop reasoning cases. Our observations reveal that critical information either fails to be properly routed to the last token position, or exhibits a weak signal, preventing effective reasoning. This explains why current KE methods underperform (§3.1): they optimize for isolated parameter changes rather than circuit-level integration needed for compositional reasoning, as shown in Figure 2 (b).
@@
To bridge this fundamental gap, we propose Circuit-aware Knowledge Editing (CaKE) in §3.2. By explicitly aligning edits with the LLM’s native reasoning architecture, CaKE transforms static knowledge updates into generalizable knowledge reasoning circuit—models that not only store static edited facts but also dynamically apply them in downstream reasoning tasks. Specifically, for the updated knowledge, we design circuit-aware tasks that require the LLM to leverage this new knowledge for latent reasoning, ensuring its integration across distinct segments of the reasoning circuit, as depicted in Figure 2(c). To prevent unintended data leakage, we construct these tasks using ad-hoc features Zhang et al. (2024c) that are temporarily associated with the entities, such as ‘Japan is colored green. The capital city of the country colored in green is’. We then guide LLMs to construct the reasoning circuits to utilize the updated knowledge by training with the curated data. Extensive experiments (§4) demonstrate the effectiveness of CaKE, outperforming existing knowledge editing methods on the multi-hop editing benchmarks MQuAKE on LLMs of different sizes, including LLAMA3-8B-Instruct Dubey et al. (2024), Qwen2.5-7B-Instruct Yang et al. (2024a), and LLAMA3-70B-Instruct.
@@ 2 Analyzing Reasoning Circuit in LLM
In this part, we integrate previous findings on multi-hop reasoning and circuit analysis Yao et al. (2024); Biran et al. (2024); Yang et al. (2024b); Wang et al. (2024e) and factual recall Merullo et al. (2024) to illustrate how language models use knowledge to tackle multi-hop reasoning tasks. Based on this, we can better understand the cause of failures for multi-hop reasoning.
@@ 2.2 Multi-hop Reasoning Circuit
Building on the insights from prior work Biran et al. (2024); Yao et al. (2024), we can define a structured circuit mechanism for multi-hop reasoning in transformer-based LLMs, as illustrated in Figure 2(a). The three distinct computational phases: 1) The model processes the initial relation $r_{1}$ and entity $e_{1}$, encoding the bridge entity $e_{2}$ in the final token position of the first prompt segment ($t_{1}$). 2) Critical features, including $e_{2}$ and the second relation, $r_{2}$ are transferred to the last token position $t_{2}$, preparing for final resolution. 3) The model computes the target $e_{3}$ by resolving $r_{2}$ and $e_{2}$, giving the result in the final token position.
Model Entity Patch Relation Patch
LLaMA3-8B-Ins. 85.35 56.20
Qwen2.5-7B-Ins. 97.29 55.40
Table 1: Activation Patching Success Rates (%).
Hence, based on the linearity theory Hernandez et al. (2024), multi-hop reasoning in LLM can be formalized as:
@@
Evaluation
To validate this circuit hypothesis, we conduct a causal analysis to determine whether modifying the variables in the function $F$ leads to corresponding changes in the output. Our intervention strategy focuses on the critical last token position for the second hop in Figure 9, where intermediate variables are stored. Specifically, we consider: 1).Entity Patching: Replacing the representation of $e_{2}$ with an alternative entity $e_{patch}$. For example, given the prompt ‘The official currency of the country where the Eiffel Tower is located is’, we substitute the representation of the bridge entity ‘France’ with ‘China’, expecting the output to change to ‘Renminbi’. 2).Relation Patching: Replacing the representation of $r_{2}$ with an alternative relation $r_{patch}$. For instance, given the same prompt, we substitute the representation of ‘official currency’ with ‘capital’, expecting the output to change to ‘Paris’. A successful patching (producing $F_{2}(e_{patch},r_{2})$ or $F_{2}(e_{2},r_{patch})$) would confirm the model’s reliance on these specific representations for reasoning. We conduct experiments on LLAMA3-8B-Instruct and Qwen2.5-7B-Instruct and employ PatchScope Ghandeharioun et al. (2024) for targeted activation patching (detailed in §B.1). Table 1 shows that in both the LLAMA3-8B and Qwen2.5-7B models, substituting variable representations at the last token position leads to corresponding behavioral changes, particularly in entity patching, where accuracy exceeds 80%. These results provide mechanistic evidence for the reasoning circuit we identified before.
@@
Table 2: The results of LLAMA3-8B-Instruct (32 layers) and Qwen2.5-7B-Instruct (28 layers). Cases are the percentage of data we can detect the information, and Layer is the mean of the first layer in which we detect the information.
@@ 2.3 Circuit in Failure Phenomena
Then, we aim to understand why language models sometimes fail at multi-hop reasoning despite successfully answering individual single-hop questions. For instance, a model may correctly answer ‘the capital of Russia’ with ‘Moscow’ and ‘the country of citizenship of Fyodor Dostoyevsky’ with ‘Russia’, yet fail to answer the multi-hop question ‘the capital of the country of citizenship of Fyodor Dostoyevsky is’ correctly. To systematically analyze this issue, we focus on the second hop of reasoning, as the model typically performs well on the first hop. We categorize the data from the HoppingTooLate dataset [^1] We filter out short-cut cases as done by Biran et al. (2024). into three subsets based on the model’s behavior: Correct: The model answers both single-hop questions $(e_{1},r_{1},e_{2})$ and $(e_{2},r_{2},e_{3})$ correctly, as well as the multi-hop question $(e_{1},r_{1},r_{2},?)$. Inconsistent: The model answers both single-hop questions correctly but fails on the multi-hop question. However, some cases in the Correct set contain the bridge entity $e_{2}$ but have a different subject $e_{1}^{\prime}$ and relation $r_{1}^{\prime}$, for which the model correctly answers $(e_{1}^{\prime},r_{1}^{\prime},r_{2},?)$. This suggests that while the model can leverage knowledge in some contexts, it fails to generalize, indicating reasoning gaps rather than missing knowledge. Incorrect: The model answers both single-hop questions correctly but fails on the multi-hop question in all contexts $(e_{1}^{\prime},r_{1}^{\prime},e_{2})$. This implies a complete failure to employ the knowledge for multi-hop reasoning. To investigate these failure modes, we check whether the models construct the reasoning circuit by monitoring key variables ($e_{1}$, $e_{2}$, and $r_{2}$) at critical positions ($t_{1}$ and $t_{2}$) across the model’s layers using the PatchScope as Biran et al. (2024) did. Our analysis reveals several interesting patterns, extending beyond the ‘hopping too late’ problem identified by Biran et al. (2024).
@@
We list the results in Table 2 (more details in Figure 8 in Appendix). For the correct subset, we observe strong evidence of the reasoning circuit functioning as expected: a large portion of $e_{2}$ is detected at both $t_{1}$ ($e_{2}$ from $t_{1}$) and $t_{2}$ ($e_{2}$ from $t_{2}$) in both LLAMA3 and Qwen2.5 models. The model correctly uses the $r_{2}$ and $e_{2}$ information at $t_{2}$ to produce the final answer $e_{3}$. Contrastly, in the Incosistent subsets, we can find that despite detecting $e_{2}$ and $r_{2}$ at $t_{2}$, the model often fails to produce the correct $e_{3}$ answer ($e_{3}$ from $t_{2}$: only 22.7% in LLAMA3 and 39.4% in Qwen2.5 of cases we can detect at $t_{2}$). We hypothesize that the $e_{2}$ information, though present, may be insufficient to trigger the second-hop reasoning circuit, leading to the failure to execute the function $F(e_{2},r_{2})$ effectively. What’s more, in the Incorrect subsets, we can find the needed $e_{2}$ information is rarely detected at the $t_{2}$ position ($e_{2}$ from $t_{2}$: Only 17.7% in LLAMA3 and 18.9% in Qwen2.5). Even when $e_{2}$ is detected, it typically emerges in much later layers (layer 21 in LLAMA3 and layer 13.5 in Qwen2.5), making it too late to be effectively utilized for the second-hop computation, aligned with Biran et al. (2024)’s findings. We conjecture the model fails to propagate $e_{2}$ to the $t_{2}$ position, resulting in the variable $e_{2}$ missing for conducting the $F(e_{2},r_{2})$ function.
@@
To test our hypothesis, we conduct interventions to enhance the information at the specific positions to see if we can improve the model’s performance in these failure cases. We test three ways: back-patching the $t_{1}$ and $t_{2}$ position as Biran et al. (2024) did, which would enhance the information at the position, and cross-position patching the information from $t_{1}$ to the $t_{2}$ position, which explicitly propagates the information from $t_{1}$ to $t_{2}$ (details in Figure 10 in Appendix). From the results in Figure 3, we can find a high success rate for all the inconsistent and incorrect cases, but they demonstrate different paradigms. For the inconsistent cases, back-patching would lead to better performance, while for the incorrect cases, patching knowledge from the $t_{1}$ to $t_{2}$ usually shows better outcomes. This proves our previous hypothesis that for the incorrect cases, due to the propagation failure, the model fails to move the $e_{2}$ to $t_{2}$ position, and manual routing via cross-patching can mitigate the issue. Meanwhile, for inconsistent cases, amplification via back-patching compensates the weak signal when valid $e_{2}$ representations reach $t_{2}$ but lack sufficient magnitude for subsequent reasoning.

### v2

@@ Abstract
Knowledge Editing (KE) enables the modification of outdated or incorrect information in large language models (LLMs). While existing KE methods can update isolated facts, they often fail to generalize these updates to multi-hop reasoning tasks that rely on the modified knowledge. Through an analysis of reasoning circuits—the neural pathways LLMs use for knowledge-based inference, we find that current layer-localized KE approaches (e.g., MEMIT, WISE), which edit only single or a few model layers, inadequately integrate updated knowledge into these reasoning pathways. To address this limitation, we present CaKE (Circuit-aware Knowledge Editing), a novel method that enhances the effective integration of updated knowledge in LLMs. By only leveraging a few curated data samples guided by our circuit-based analysis, CaKE stimulates the model to develop appropriate reasoning circuits for newly incorporated knowledge. Experiments show that CaKE enables more accurate and consistent use of edited knowledge across related reasoning tasks, achieving an average improvement of 20% in multi-hop reasoning accuracy on the MQuAKE dataset while requiring less memory than existing KE methods. We release the code and data in https://github.com/zjunlp/CaKE.
@@
[…]
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in diverse tasks Yang et al. (2024a); Azaria et al. (2024); Dubey et al. (2024); OpenAI (2024); Guo et al. (2025), achieving performance that rivals or even exceeds human experts. However, their practical deployment faces some critical limitations: parametric knowledge remains static after pretraining, making it challenging to keep up with evolving real-world information; their propensity for hallucinations also undermines reliability Chen (2024). Knowledge editing (KE) has emerged as a promising solution to update the knowledge in models precisely Mitchell et al. (2021); Wang et al. (2024c); Jiang et al. (2025). Although existing KE methods achieve good results on simple factual updates Yao et al. (2023); Zhang et al. (2024b), they often exhibit fundamental limitations: edits propagate inconsistently through related knowledge structures and downstream reasoning tasks Cohen et al. (2024); Qin et al. (2024); Yao et al. (2023); excessive focus on surface-level pattern matching Hoelscher-Obermaier et al. (2023), and locality issues for other unrelated knowledge and general ability Gu et al. (2024); Gupta et al. (2024).
@@
Our work specifically addresses the poor performance of edited models in downstream reasoning tasks that involve the updated knowledge Zhong et al. (2023); Zhang et al. (2024d). Consider a representative case in Figure 1 : after editing ‘Eddie Mathews, citizenship, United States → United Kingdom’, models correctly answer direct queries but fail multi-hop reasoning like ‘The capital of the country that Eddie Mathews was a citizen of is?’ (still outputting ‘Washington D.C.’). Critically, this is not merely an editing artifact: vanilla LLMs often correctly answer single-hop questions while failing their multi-hop counterparts Yang et al. (2024b); Biran et al. (2024), suggesting deeper architectural limitations in knowledge utilization.
@@
We trace these limitations to a misalignment between KE strategies and the inherent reasoning architectures of LLMs. To investigate this disconnect, we examine how LLMs leverage knowledge in downstream reasoning tasks. Recent analysis suggests that knowledge is not merely statically stored but dynamically activated through specialized circuits Yao et al. (2024); Biran et al. (2024); Yu et al. (2025). However, these analyses overlook the phenomenon of LLM failures in reasoning circuits and fail to explore the underlying causes. Our investigation (§2) delves deeper into reasoning circuits, analyzing their structure and identifying the reasons behind failures in multi-hop reasoning. Specifically, the multi-hop reasoning emerges from coordinated computing circuits: early layers handle the first hop, extracting the bridge entity at the end-token of the first hop. This bridge entity, along with second-hop relation information, is then routed to the last token position in the middle layers. Subsequently, later layers utilize this information at the last token position to complete the reasoning process (Figure 2 (a)). We then analyze the entity and relation information at the last token position in failed multi-hop reasoning cases. Our observations reveal that critical information either fails to be properly routed to the last token position, or exhibits a weak signal, preventing effective reasoning. This explains why current KE methods underperform (§3.1): they optimize for isolated parameter changes rather than circuit-level integration needed for compositional reasoning (Figure 2b).
@@
To bridge this fundamental gap, we propose Circuit-aware Knowledge Editing (CaKE) in §3.2. Unlike methods that only update localized static knowledge, CaKE actively constructs reasoning circuits that enable dynamic application of edited knowledge in downstream tasks. We first design circuit-aware training data that integrates across distinct segments of the reasoning process to force the LLM to leverage updated knowledge for latent reasoning (Figure 2c). Remarkably, we find that only a few such samples are sufficient to integrate knowledge across the reasoning circuit while maintaining strong general ability. Moreover, to prevent unintended data leakage, we construct these data using ad-hoc features Zhang et al. (2024c) that are temporarily associated with the entities, such as ‘Japan is colored green. The capital city of the country colored in green is’. Finally, we guide LLMs to establish reasoning circuits by training with the curated data. Extensive experiments (§4) demonstrate CaKE’s effectiveness: it outperforms existing knowledge editing methods on the MQuAKE multi-hop benchmark for both LLAMA3-8B-Instruct Dubey et al. (2024) and Qwen2.5-7B-Instruct Yang et al. (2024a). Notably, CaKE achieves this while being more memory-efficient than alternatives and successfully scales to larger models like LLAMA3-70B-Instruct.
@@ 2.2 Multi-hop Reasoning Circuit
Building on the insights from prior work Biran et al. (2024); Yao et al. (2024), we can define a structured circuit mechanism for multi-hop reasoning in transformer-based LLMs, as illustrated in Figure 2(a). The three distinct computational phases: 1) The model processes the initial relation $r_{1}$ and entity $e_{1}$, encoding the bridge entity $e_{2}$ in the final token position of the first prompt segment ($t_{1}$). 2) Critical features, including $e_{2}$ and the second relation, $r_{2}$ are transferred to the last token position $t_{2}$, preparing for final resolution. 3) The model computes the target $e_{3}$ by resolving $r_{2}$ and $e_{2}$, giving the result in the final token position. Hence, based on the linearity theory Hernandez et al. (2024), multi-hop reasoning in LLM can be formalized as:
@@
Table 1: The results of LLAMA3-8B-Instruct (32 layers) and Qwen2.5-7B-Instruct (28 layers). Cases are the percentage of data we can detect the information, and Layer is the mean of the earliest layer where the required information is detected.
@@ 2.3 Circuit in Failure Phenomena
Then, we aim to understand why language models sometimes fail at multi-hop reasoning despite successfully answering individual single-hop questions. For instance, a model may correctly answer ‘the capital of Russia’ with ‘Moscow’ and ‘the country of citizenship of Fyodor Dostoyevsky’ with ‘Russia’, yet fail to answer the multi-hop question ‘the capital of the country of citizenship of Fyodor Dostoyevsky is’ correctly. To systematically analyze this issue, we focus on the second hop of reasoning, as the model typically performs well on the first hop. We categorize the data from the HoppingTooLate dataset [^1] We filter out short-cut cases as done by Biran et al. (2024). into three subsets based on the model’s behavior: Correct: The model answers both single-hop questions $(e_{1},r_{1},e_{2})$ and $(e_{2},r_{2},e_{3})$ correctly, as well as the multi-hop question $(e_{1},r_{1},r_{2},?)$. Inconsistent: The model answers both single-hop questions correctly but fails on the multi-hop question. However, we observe that some questions in the Correct set share the same ’bridge’ entity $e_{2}$, even though they originate from distinct subject-relation pairs, that the model answers correctly. $(e_{1}^{\prime},r_{1}^{\prime},r_{2},?)$. This suggests that while the model can leverage knowledge in some contexts, it fails to generalize, indicating reasoning gaps rather than missing knowledge. Incorrect: The model answers both single-hop questions correctly but fails on the multi-hop question in all contexts $(e_{1}^{\prime},r_{1}^{\prime},e_{2})$. This implies a complete failure to employ the knowledge for multi-hop reasoning. To investigate these failure modes, we check whether the models construct the reasoning circuit by monitoring key variables ($e_{1}$, $e_{2}$, and $r_{2}$) at critical positions ($t_{1}$ and $t_{2}$) across the model’s layers using the PatchScope as Biran et al. (2024) did. Our analysis reveals several interesting patterns, extending beyond the ‘hopping too late’ problem identified by Biran et al. (2024).
@@
We list the results in Table 1. For the correct subset, we observe strong evidence of the reasoning circuit functioning as expected: a large portion of $e_{2}$ is detected at both $t_{1}$ ($e_{2}$ from $t_{1}$) and $t_{2}$ ($e_{2}$ from $t_{2}$) in both LLAMA3 and Qwen2.5 models. The model correctly uses the $r_{2}$ and $e_{2}$ information at $t_{2}$ to produce the final answer $e_{3}$. Contrastly, in the Incosistent subsets, we can find that despite detecting $e_{2}$ and $r_{2}$ at $t_{2}$, the model often fails to produce the correct $e_{3}$ answer ($e_{3}$ from $t_{2}$: only 22.7% in LLAMA3 and 39.4% in Qwen2.5 of cases we can detect at $t_{2}$). We hypothesize that the $e_{2}$ information, though present, may be insufficient to trigger the second-hop reasoning circuit, leading to the failure to execute the function $F(e_{2},r_{2})$ effectively. What’s more, in the Incorrect subsets, we can find that the needed $e_{2}$ information is rarely detected at the $t_{2}$ position ($e_{2}$ from $t_{2}$: Only 17.7% in LLAMA3 and 18.9% in Qwen2.5). Even when $e_{2}$ is detected, it typically emerges in much later layers (layer 21 in LLAMA3 and layer 13.5 in Qwen2.5), making it too late to be effectively utilized for the second-hop computation, aligned with Biran et al. (2024)’s findings. We conjecture the model fails to propagate $e_{2}$ to the $t_{2}$ position, resulting in the variable $e_{2}$ missing for conducting the $F(e_{2},r_{2})$ function.
@@
To test our hypothesis, we conduct interventions to enhance the information at the detected layers to see if we can improve the model’s performance in these failure cases. We test three ways: back-patching the $t_{1}$ and $t_{2}$ position as Biran et al. (2024) did, which would enhance the information at the position, and cross-position patching the information from $t_{1}$ to the $t_{2}$ position, which explicitly propagates the information from $t_{1}$ to $t_{2}$ (details in Figure 7 in Appendix). From the results in Figure 3, we can find a high success rate for all the inconsistent and incorrect cases, but they demonstrate different paradigms. For the inconsistent cases, back-patching would lead to better performance, while for the incorrect cases, patching knowledge from the $t_{1}$ to $t_{2}$ usually shows better outcomes. This proves our previous hypothesis that for the incorrect cases, due to the propagation failure, the model fails to move the $e_{2}$ to $t_{2}$ position, and manual routing via cross-patching can mitigate the issue. Meanwhile, for inconsistent cases, amplification via back-patching compensates the weak signal when valid $e_{2}$ representations reach $t_{2}$ but lack sufficient magnitude for subsequent reasoning.
@@ 3 Circuits-aware Knowledge Editing
Building on our previous reasoning analysis, we rethink the reason why current knowledge editing methods fail under multi-hop reasoning circumstances despite their great performance under single-fact editing.